Shift report 20000817 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Thu Aug 17 2000 - 16:55:44 EDT

  • Next message: Peter H. L. Christiansen: "Re: Revised tpc monitor"

    ------------------------------------------------------------
    Date: Thu Aug 17 16:55:44 2000, Shift: 08:00-16:00
    Supervisor: sjs, with: sjs,b1
    Log book#4 pages 111 - 112
    
    Shift Summary:
    ------------------------------------------------------------
    1.) Beam off most of morning.  Store from 12:40 to 14:30,
    when beam dumped.
    2.) KO fixed DAQ
        Run 2322 with 28657 collision (MRS 90 deg, -4kG)
        Run 2323 Camac pedestal
        Run 2324 tpc pedestal
        D1 and D2 polarity switched (now D1 520A(B polarity),
          D2 410A (B polarity)
        Loose cables for two Si detectors fixed, several
          Si biases increased
    3.) T3 anode trips at low voltage, set off.
        BBL 16 and 17 tripped (reset)
       
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Thu Aug 17 2000 - 16:59:23 EDT